OpenStudy (anonymous):

20% of what number is 47

OpenStudy (anonymous):

\[\frac{ 20 }{ 100 }* x = 47\] \[\frac{ 20x }{ 100 } = 47\] \[20x = 47*100\] \[20x =4700\] \[x = 4700/20 = 235\] So the answer is 235.